# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ H.2 Experimental evaluation of GeoNGNN-C
We evaluate the ability of GeoNGNN-C to distinguish enantiomers and learn meaningful chirality-relevant representations through three tasks on two datasets [1, 19]: 1) Classification of tetrahedral chiral centers as R/S. R/S offers a fundamental measure of molecular chirality, and this classification task serves as an initial evaluation of the model’s capacity to discriminate enantiomers. The underlying dataset contains a total of 466K conformers of 78K enantiomers with a single tetrahedral chiral center. 2) Enantiomer ranking task and binding affinity prediction task. The two tasks share the same underlying dataset, while the second task requires models to predict the binding affinity value for each enantiomer directly, and the first requires the model to predict which enantiomer between the corresponding pair exhibits higher such values. Chirality is essential here because enantiomers often exhibit distinct behaviors when docking in a chiral protein pocket, thus leading to differences in binding affinity. The underlying dataset contains 335K conformers from 69K enantiomers carefully selected and labeled by Adams et al. [1].
@@
In the task of R/S classification, GeoNGNN-C adopts a three-layer outer GNN positioned before the inner GNNs. This configuration is designed to enhance the node features first, facilitating the subsequent extraction of chirality-related properties within the inner GNNs. For the other two tasks, GeoNGNN-C consists of five inner layers and one outer layer, with the outer GNN placed after the inner GNN. The model is trained with L1 loss. Specifically, if predicted affinity difference between two enantiomers falls below the threshold of 0.001, it is considered that the model lacks the capability to discriminate between the two enantiomers. Note that all the evaluation criteria are consistent with previous works [1, 19] for fair comparisons.
@@
We compare GeoNGNN-C against previous high-performance models, SphereNet [39], ChiRo [1], ChiENN [19], and other baseline models in [44, 1, 19]. It is worth noting that, for the last two tasks, the docking scores are labeled at the stereoisomer level, meaning that all different conformers of a given enantiomer are assigned the same label, specifically the best score, by Adams et al. [1]. As a result, 2D Graph Neural Networks (GNNs) such as DMPNN+tags and models that possess invariance to conformer-level transformations, such as bond rotations, including ChiRo [1], Tetra-DMPNN [44] and ChiENN [19], inherently exhibit significantly better data efficiency and usually exhibit better performance. GeoNGNN-C and SphereNet do not exhibit such invariance, and are trained on 5 conformers for each enantiomer as data augmentation.

### v3

@@ H.2 Experimental Evaluation of GeoNGNN-C
We evaluate the ability of GeoNGNN-C to distinguish enantiomers and learn meaningful chirality-relevant representations through three tasks on two datasets (Adams et al., 2021; Gaiński et al., 2023): 1) Classification of tetrahedral chiral centers as R/S. R/S offers a fundamental measure of molecular chirality, and this classification task serves as an initial evaluation of the model’s capacity to discriminate enantiomers. The underlying dataset contains a total of 466K conformers of 78K enantiomers with a single tetrahedral chiral center. 2) Enantiomer ranking task and binding affinity prediction task. The two tasks share the same underlying dataset, while the second task requires models to predict the binding affinity value for each enantiomer directly, and the first requires the model to predict which enantiomer between the corresponding pair exhibits higher such values. Chirality is essential here because enantiomers often exhibit distinct behaviors when docking in a chiral protein pocket, thus leading to differences in binding affinity. The underlying dataset contains 335K conformers from 69K enantiomers carefully selected and labeled by Adams et al. (2021).
@@
In the task of R/S classification, GeoNGNN-C adopts a three-layer outer GNN positioned before the inner GNNs. This configuration is designed to enhance the node features first, facilitating the subsequent extraction of chirality-related properties within the inner GNNs. For the other two tasks, GeoNGNN-C consists of five inner layers and one outer layer, with the outer GNN placed after the inner GNN. The model is trained with L1 loss. Specifically, if predicted affinity difference between two enantiomers falls below the threshold of 0.001, it is considered that the model lacks the capability to discriminate between the two enantiomers. Note that all the evaluation criteria are consistent with previous works Adams et al. (2021); Gaiński et al. (2023) for fair comparisons.
@@
We compare GeoNGNN-C against previous high-performance models, SphereNet (Liu et al., 2021), ChiRo (Adams et al., 2021), ChiENN (Gaiński et al., 2023), and other baseline models in (Pattanaik et al., 2020; Adams et al., 2021; Gaiński et al., 2023). It is worth noting that, for the last two tasks, the docking scores are labeled at the stereoisomer level, meaning that all different conformers of a given enantiomer are assigned the same label, specifically the best score, by Adams et al. (2021). As a result, 2D Graph Neural Networks (GNNs) such as DMPNN+tags and models that possess invariance to conformer-level transformations, such as bond rotations, including ChiRo (Adams et al., 2021), Tetra-DMPNN (Pattanaik et al., 2020) and ChiENN (Gaiński et al., 2023), inherently exhibit significantly better data efficiency and usually exhibit better performance. GeoNGNN-C and SphereNet do not exhibit such invariance, and are trained on 5 conformers for each enantiomer as data augmentation.